The time between requests to a web server is exponentially distributed with mean 0.5 seconds. NOTE: This is a multi-part question. Once an answer is submitted, you will be unable to return to this part.

Find the probability that the time between requests is between 1 and 2 seconds.

Probability =
Answer #1

Solution:

X : time between requests to a web server

X has exponential distribution

mean = 1/λ = 0.5 seconds. Hence λ = 2

probability that the time between requests is between 1 and 2 seconds = P( 1 < x < 2 )

  = P(X<2) - P(X<1)

=(1−e−2/0.5) - (1−e−1/0.5)

=0.9817 - 0.8647 = 0.1170

Probability = 0.1170